OpenStudy (anonymous):

can some please helo me on algebra?(:

OpenStudy (mertsj):

can you be more specific?

OpenStudy (anonymous):

does the data in the table represent a direct variation or an inverse variation? 4. x 2 4 6 8 y 14 28 42 56 (1 point)

OpenStudy (mertsj):

If it is, you always multiply x by the same number to get y

OpenStudy (anonymous):

# direct variation; y = 1/7x # inverse variation; xy = 7 # direct variation; y = 7x # inverse variation; xy = 1/7

OpenStudy (anonymous):

that was the rest of the question can you help me more please

OpenStudy (mertsj):

If x = 2, y = 14 according to the table. What do you do to 2 to get 14?

OpenStudy (anonymous):

you will need to multiply by 7 to get 14 would if be direct or inverse?

OpenStudy (mertsj):

Now look at the next two values: x = 4 and y = 28 Is it the same?

OpenStudy (anonymous):

yes it will be the same so would it be inverse xy=7